OpenStudy (anonymous):

how do you solve this equation? (x+2)^(-1/6) -5= -7? Please I have my midterm tomorrow!!

OpenStudy (anonymous):

is that written correctly ...^(-1/6) ?

OpenStudy (anonymous):

yup, the exponent is -1/6

OpenStudy (anonymous):

the teacher gave us an answer packet and said that this answer is -127/64 but I don't understand how she got that

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

Use PEMDAS

OpenStudy (anonymous):

I understand that you have to use PEMDAS, you do for most math equations...

OpenStudy (anonymous):

I kept getting -62 for some reason

OpenStudy (anonymous):

Remove negative exponents in the expression 1/(x+2)^1/6 1/(x+2)^1/6 -5 = -7 -5 does not contain the variable to solve so move it to the right hand side of the equation by adding 5 to both sides Subtract 7 from 5 to get -2 1/(x+2)^1/6 = -2 the variable is in the denominator so solve as a ratio 1/-2=(x+2)^1/6 switch the sides so X is on the left-hand side of the equation (x+2)^1/6=1/-2 Move the minus sign to the front of the expression Raise each side of the equation to the 6 power to eliminate the fractional exponent ((x+2)^1/6)^6=(-1/2)^6 Now solve for X X=-127/64

OpenStudy (anonymous):

hope that helped

OpenStudy (anonymous):

thank you so much!!!!!

OpenStudy (anonymous):

ur welcome are you good at geo?

OpenStudy (anonymous):

im okay

OpenStudy (anonymous):

do you have a question?